OpenStudy (anonymous):

Ok I am a little stuck. Let x^3+y^3=65. Find y"(x) at the point (4,1). I found Y(') but cant seem to get the answer for the next double prime.

OpenStudy (anonymous):

I think y(') is \[\frac{ -x^2 }{ y^2 }\]

OpenStudy (cruffo):

I got that as well.

OpenStudy (cruffo):

Are you familiar with quotient rule?

OpenStudy (anonymous):

Yea thats how I set it up next but I must have made a mistake

OpenStudy (anonymous):

I had (-2xy^2+x^2(2y)dy/dx)/y^4

OpenStudy (cruffo):

On the top, when you take derivative of y^2 should get 2yy' :)

OpenStudy (anonymous):

oh thats what I meant where the dy/dx is

OpenStudy (cruffo):

man, sorry. That's right!

OpenStudy (anonymous):

I dont use y' b/c I always make a mistake lol

OpenStudy (anonymous):

I think my problem is in simple algebra that I can get the y' alone :(

OpenStudy (cruffo):

don't need to. Substitute \( \large \dfrac{dy}{dx} = \dfrac{-x^2}{y^2}\)

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

I forgot part of the problem is says y"(4)= is what its asking

OpenStudy (cruffo):

Oh, good. So you just need to substitute x = 4, y = 1, and y' = -16

OpenStudy (anonymous):

where did u get 16?

OpenStudy (cruffo):

y'(4) = -(4)^2/(1)^2

OpenStudy (anonymous):

ahh

OpenStudy (cruffo):

y''(4) = -520?

OpenStudy (anonymous):

Yeah ty!

OpenStudy (cruffo):

np!

OpenStudy (anonymous):

I only have 4 more left :P this homework is so long lol